OpenStudy (anonymous):

If 10.0 mL of phosphoric acid solution required 23.5 mL of .200 M sodium hydroxide solution for neutralization to a phenolphthalein end point, calculate the molarity of the phosphoric acid. H3PO4(aq)+3 NaOH (aq) --> Na3PO(aq)+3H2O

OpenStudy (somy):

Step 1 use formula molarity= mole/volume molarity= mole/L you have molarity and volume (convert to L) for NaOH thus find mole of NaOH

OpenStudy (somy):

Step 2 as you find mole of NaOH work with ratio what is the ratio between NaOH and H3PO4? ratio can be found my looking at coefficients ( numbers before compounds, if there is no number written, that means its 1) \[\frac{ mole ~of~ NaOH }{ coefficient~ of~ NaOH } =\frac{ mole ~of~ H_3PO_4 }{ coefficient ~ of ~H_3PO_4 }\] thus find mole of H3PO4

OpenStudy (somy):

Step 3 you found mole of Phosphoric acid and you have volume given to you (don't forget to convert to Liters) so use formula molarity = mole/volume and get the final answer Hope thats helpful :)

OpenStudy (anonymous):

It did thank you!!!

OpenStudy (somy):

you are welcome :))